An investment had a nominal return of 10.4 percent last year. If the real return on the investment was only 5.9 percent, What was the inflation rate for the year?
Answer #1

Nominal Rate =10.4%
Real Rate of Return =5.9%
Inflation Rate =(1+Nominal Rate)/(1+Real Rate of Return)-1 =(1+10.4%)/(1+5.9%)-1 =4.25%